Fermi National Laboratory
 

Operation Summary for July 13, 2011 - July 15, 2011

 

Wednesday July 13

The day shift began with NuMI off due to target replacement, with Operations monitoring the Tevatron (TEV) store 8884 (111.84E30) and the Recycler stash (363.2E10), with the Antiproton Source (Pbar) stacking (7.9E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, operators prepared Meson enclosures M01-5 for access at 8:16 AM. Water technicians continued to work on TEV heat exchangers. Operations terminated store 8884 at 10:47 AM. At 11:03 AM, operators reported that a heater firing unit (HFU) at TEV sector F2 fired; operators dumped the TEV ramp so that techs could replace the unit. Operators prepared D0 for a one-hour access at 11:12 AM. Cryo System techs reported at 12:20 PM that they had finished their work at E3 and A4. Operations put the TEV through a dry squeeze at 1:09 PM and a wet squeeze at 1:46 PM. Operations put the TEV into shot setup at 2:42 PM and established store 8888 at 3:50 PM with an average initial luminosity (AIL) of 389.77E30. At 3:58 PM, operators reported that the TEV E2 air compressor tripped off; operators reset it. Operators reported at the end of the shift that Booster personnel were conducting RF repairs.

During the evening shift, operators reported at 4:22 PM that they had prepared MI for access to check for LCW leaks. Booster personnel reported at 4:45 PM that they had completed their RF work. People who were searching the MI tunnel for LCW leaks reported at 4:52 PM that they could not find any leaks. The Cryo Coordinator reported at 4:57 PM that there was a problem with the A4 wet engine that would have to be repaired in between stores. Operations established beam to MiniBooNE at 6:05 PM. At 6:27 PM, operators reported that they were having trouble closing the MI-10 Lower VCB; operators talked with an EE Support person and then were able to close the VCB. Pbar resumed stacking at 6:38 PM. Operators reported short downtimes due to a Linac RF station (LRF1).

Thursday July 14

The midnight shift began with Operations monitoring store 8888 (113.61E30) and the stash (227.6E10), with Pbar stacking (17.1E10), and with MiniBooNE taking beam.

During the midnight shift, operators reported at 2:34 AM that a MI RF station (MIRF4) tripped off on a LCW flow indication and would not reset; operators bypassed the station. At 3:07 AM, the MI South Anode Power Supply tripped off; it reset. At 3:36 AM, the MI-40 LCW pump #2 tripped off and would not reset. Pbar personnel reported at 6:52 AM that they had walked through AP0, AP10. AP30, and AP50 in search of their LCW leak, but couldn’t find anything. Operations terminated store 8888 at 7:33 AM. Cryo System technicians began working on the TEV sector A4 wet engine at 7:43 AM. Operators prepared D0 for access at 7:48 AM.

During the day shift, Cryo System techs reported at 8:17 AM that they had completed their work on the A4 wet engine. D0 came out of access at 8:31 AM. Operators reported that they had put the TEV through a dry squeeze and a wet squeeze by 9:46 AM. Operations put the TEV into shot setup at 10:10 AM. An RF person reported finding the MI LCW leak coming from the MI Center Anode Power Supply, in the MI-60 service building. Operations established store 8891 at 11:20 AM with an AIL of 355.96E30. At 11:55 AM, store 8891 aborted when the TEV sectors D1 and D2 vacuum valve positions went to an indeterminate state. A Vacuum person reported that the problem was in the tunnel. Operations established beam to MTA at 12:09 PM. Operators prepared the TEV for access at 12:10 PM. Operators reported that the TEV D3 VCB would not open. Operators prepared D0 for an access at 12:33 PM. EE Support personnel investigated the TEV D3 VCB and got it to open at 12:53 PM. A Water Works person reported that the problem with the MI-40 LCW pump #2 was a bad contactor. MI personnel began a coalescing study at 12:57 PM. Cryo system technicians began work on the TEV C4 wet engine; they completed the C4 work about two hours latter. Operations disabled beam to MTA at 1:41 PM. Operations disabled the MiniBooNE beam at 1:57 PM, due to the need to access the MI-12 service building for HVAC work. Vacuum personnel came out of the TEV at 2:26 PM; they reported replacing a solenoid on a beam valve (D2BVU). Operators, while turning on the TEV, reported a ground fault. Operators and the TEV Operations Specialist investigated.

During the evening shift, during the process of trying to narrow down the location of the TEV ground fault, the fault disappeared. At 7:24 PM, operators reported problems with the TEV Bypass Loop; after a lot of work by a lot of people, the problem was found while running through the TEV QPM and Power Supply checkout: it was a bad setting for both the high and low voltage to ground limits. At 8:42 PM, a Linac Klystron RF station (KRF6) tripped off and wouldn’t reset; Linac personnel came in and discovered, at 11:49 PM, a LCW leak inside the charging power supply for KRF6; they began repairs.

Friday July 15

The midnight shift began with no Linac beam due to KRF6 work and with Operations monitoring the stash (322.6E10) and the Pbar stack (3.8E10).

During the midnight shift, at 12:09 AM, LRF4 tripped off and wouldn’t reset. Linac personnel reported at 3:32 AM that they had completed their work on KRF6 and LRF4. Linac resumed sending beam at 3:39 AM. Operations put the TEV into shot setup at 5:08 AM and established store 8894 at 6:17 AM with an AIL of 356.95E30. At 7:35 AM, BRF12 tripped off for the first of four times in ten minutes; it reset.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow the following events today: Pbar studies, Booster studies, MI coalescing studies, MTA to take beam, and for Mechanical personnel to make up water to the MI LCW system.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/15/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory